# Large Language Model Evaluation via Matrix Nuclear-Norm

### v1

@@ 5.1.2 Language Models
5.1.3 Language Datasets
Pretraining Datasets. All sizes of the Cerebras-GPT models were pretrained on the Pile dataset. Due to resource constraints, we randomly selected a subset of 10,000 data samples from Wikipedia (Foundation, 2024) and openwebtext2 (Skylion007, 2019) to observe the Matrix Nuclear-Norm, as these datasets are part of the Pile. In addition to the datasets used for model pretraining, we incorporated supplementary datasets that were not directly involved in the training process.
Instruction-Tuning Datasets. To assess the effectiveness of the Matrix Nuclear-Norm, we selected the dolly-15k dataset (Conover et al., 2023), generated by human workers, as one of our instruction datasets. We specifically focused on the "context" portion of this dataset, which contains more informative text. This dataset will also be referenced in Section 5.3.2 for analyzing context length, prompt learning, and contextual operations.
RLHF Dataset. We utilized the hh-rlhf dataset (Bai et al., 2022), which contains human preference data related to usefulness and harmlessness, as our Reinforcement Learning with Human Feedback (RLHF) dataset. Each entry includes a pair of texts, one marked as "chosen" and the other as "rejected." We input the "chosen" portions of the dataset into the model and subsequently calculated the performance of the Matrix Nuclear-Norm.
Benchmark Datasets. To evaluate the performance of the Matrix Nuclear-Norm, we selected three benchmark datasets: OpenBookQA (Mihaylov et al., 2018), Winogrande (Sakaguchi et al., 2021), and PIQA (Bisk et al., 2020). These benchmarks are structured in a multiple-choice format. We used each question along with its correct answer as input to compute the respective evaluation metrics.
Prompt Learning Datasets. We conducted prompt learning experiments using the OpenOrca dataset (Lian et al., 2023b), where each question-answer (QA) pair features a unique system prompt. To explore the impact of these prompts on Matrix Nuclear-Norm, we carefully selected three generally instructive prompts (see Table 13) to observe their roles in the QA pairs.
Validation Datasets. Additionally, we performed inference and evaluation on two widely used datasets: AlpacaEval (Dubois et al., 2024) and Chatbot Arena (Chiang et al., 2024). We conducted inference on both datasets, evaluated the responses using our proposed metrics, and ranked the models accordingly.
@@ 5.2 Matrix Nuclear-Norm Observation
5.2.1 A Comparative Analysis of Computational Time
@@
The significant efficiency gain is attributed to the lower computational complexity of Matrix Nuclear-Norm, $O(m\cdot n+n\log n)$, compared to Matrix Entropy’s $O(n^{3})$. This makes Matrix Nuclear-Norm not only an effective but also a highly efficient metric for evaluating LLMs, especially as models continue to scale in size.
@@
Model Size Matrix Entropy Time (s) Matrix Nuclear-Norm Time (s) Ratio
 [^1] M 623.5367 72.6734 8.5800
256M 1213.0604 110.8692 10.9414
590M 2959.6949 184.7785 16.0175
1.3B 6760.1893 379.0093 17.8365
2.7B 12083.7105 732.6385 16.4934
6.7B 38791.2035 1598.4151 24.2685
13B 59028.4483 2984.1529 19.7806
Table 1: CEREBRAS-GPT: Time Comparison between Matrix Entropy and Matrix Nuclear-Norm
@@ 5.2.2 Scaling Law of Matrix Nuclear-Norm
To affirm Matrix Nuclear-Norm’s efficacy as an evaluative metric, we evaluated Cerebras-GPT models on four datasets including dolly-15k, Wikipedia, openwebtext2, and hh-rlhf comparing Matrix Nuclear-Norm, matrix entropy, perplexity, and loss. Results, detailed in Table 11 (Appendix), demonstrate Matrix Nuclear-Norm’s consistent decrease with model size enlargement, signifying better data compression and information processing in larger models. This trend (see in Figure 2(d)) validates Matrix Nuclear-Norm’s utility across the evaluated datasets. Notably, anomalies at the 2.7B and 13B highlight areas needing further exploration.
(a) Cross-Entropy Loss
(b) Perplexity
(c) Matrix Entropy
(d) Matrix Nuclear-Norm
Figure 2: Comparison of Matrix Nuclear-Norm, matrix entropy, loss, and perplexity when model scales up.
@@ 5.3.1 Sentence Operation Experiments
Figure 3 clearly indicates that sentence manipulations significantly influence Matrix Nuclear-Norm values, which generally decline as model size increases. This trend confirms the enhanced information compression capabilities of larger models. The ranking of Matrix Nuclear-Norm values by operation is as follows: Reverse $>$ Shuffle & Reverse $>$ Shuffle $>$ Base. This indicates that disrupting sentence structure through Reverse and Shuffle & Reverse operations leads to higher Matrix Nuclear-Norm values due to increased information chaos and processing complexity. In contrast, the Shuffle operation has minimal effect on compression, while the Base condition consistently yields the lowest Matrix Nuclear-Norm values, signifying optimal information compression efficiency with unaltered sentences.
Despite the overall downward trend in Matrix Nuclear-Norm values with increasing model size, the 2.7B model exhibits slightly higher values for Shuffle and Base operations compared to the 1.3B model. This anomaly suggests that the 2.7B model may retain more nuanced information when handling shuffled data or operate through more intricate mechanisms. However, this does not detract from the overarching conclusion that larger models excel at compressing information, thereby demonstrating superior processing capabilities.
@@ 5.3.2 Analysis of Length Dynamics
Figure 4: The Matrix Nuclear-Norm values for contexts of varying lengths show that as text length increases, the Matrix Nuclear-Norm continues to rise and tends to converge.
The analysis reveals that Matrix Nuclear-Norm values generally increase as input length rises, aligning with our expectations (see Figure 4). Longer inputs necessitate that the model manage and compress more information, which naturally leads to higher Matrix Nuclear-Norm values. Most models exhibit this trend, indicating effective handling of the increased information load.
However, the gpt-2.7B and gpt-13B models display anomalies in their Matrix Nuclear-Norm values at 64 and 128 tokens, where the value at 128 tokens is lower than that at 64 tokens. This discrepancy may be attributed to these models employing different information compression mechanisms or optimization strategies tailored to specific input lengths, allowing for more effective compression at those lengths.
Overall, aside from a few outliers, the results largely conform to expectations, demonstrating that Matrix Nuclear-Norm values increase with input length, reflecting the greater volume and complexity of information that models must handle.
To address the observed trend of rising Matrix Nuclear-Norm values with longer sentences, we incorporated a normalization step in our methodology via dividing the Matrix Nuclear-Norm values by the sentence length. This adjustment helps mitigate any biases introduced by models that tend to generate longer sentences during inference.
GPT MODEL SIZE
LENGTH  [^1] M 256M 590M 1.3B 2.7B 6.7B 13B
64 0.4574 0.4125 0.3787 0.3486 0.4053 0.3315 0.4148
128 0.5293 0.4680 0.4270 0.3835 0.4143 0.3477 0.4032
512 0.7883 0.6978 0.6251 0.5554 0.5265 0.4468 0.4422
1024 0.9132 0.8787 0.7802 0.6953 0.6351 0.5383 0.5028
Table 2: Analysis of Length Dynamics
@@ 5.3.3 Analysis of Prompt Learning
The experimental results (shown in Table 3) indicate that we performed inference on different sizes of GPT models using three carefully selected prompts (shown in Table 13) and calculated the Matrix Nuclear-Norm values of their responses. As the model size increased, the Matrix Nuclear-Norm values gradually decreased, demonstrating that larger models possess greater information compression capabilities. The prompts significantly influenced Matrix Nuclear-Norm, with variations reflecting the models’ responses to prompt complexity. Specifically, GPT-1.3B showed a notable decrease in Matrix Nuclear-Norm after the input prompts, indicating its sensitivity to them, while GPT-2.7B exhibited smaller changes. In contrast, GPT-6.7B displayed minimal variation across all prompts, suggesting stable performance regardless of prompt detail. Overall, more detailed prompts resulted in larger information volumes in the model’s responses, leading to corresponding changes in Matrix Nuclear-Norm values.
Table 3: Results of prompt learning with (Empty Prompt) and without (Prompt 1, 2, 3) the use of prompts. Incorporating prompts as prefixes before the QA pairs enhances the models’ ability to achieve better compression.
@@
6 Implementing Proposed Metrics: Evaluating and Ranking Language Models in Practice
@@ 6.1 Inference-Based Model Assessment
For instance, the Llama-3 70B model demonstrated superior compression capabilities compared to its 8B counterpart, as reflected by significantly lower Matrix Nuclear-Norm values across both benchmarks (see Table A.2 in the Appendix). A similar trend was observed in the Vicuna family, where Matrix Nuclear-Norm values consistently decreased from 0.4623 for the 7B model to 0.3643 for the 33B model on the AlpacaEval dataset, indicating progressive improvements in information handling (see Table 4). Additionally, the DeepSeek models exhibited a consistent decrease in Matrix Nuclear-Norm values as model size increased, further demonstrating the metric’s validity.
@@
Model DataSet 7B 13B 33B
@@
6.2 Matrix Nuclear-Norm Benchmarking: Ranking Mid-Sized Models
@@
MODEL Matrix Nuclear-Norm Rank
@@
DeepSeek-7B 0.3352 0.4357 0.3855 $\downarrow$
Gemma-7B 0.3759 0.3998 0.3879 $\downarrow$
Vicuna-7B 0.4623 0.4824 0.4724 $\downarrow$
LLaMA 2-7B 0.4648 0.5038 0.4843 $\downarrow$
QWEN 1.5-7B 0.4866 0.5165 0.5016 $\downarrow$
Mistral-7B 0.4980 0.5126 0.5053 $\downarrow$
QWEN 2-7B 0.5989 0.5751 0.5870 $\downarrow$
QWEN 1.5-72B 0.5291 0.5065 0.5178 $\downarrow$
QWEN 2-72B 0.5261 0.4689 0.4975 $\downarrow$
Llama 3-70B 0.4935 0.4967 0.4951 $\downarrow$
Llama 2-70B 0.3862 0.4086 0.3974 $\downarrow$
Table 5: Matrix Nuclear-Norm Rankings: A Comparative Analysis of Model Performance
To validate the design rationale and robustness of the Matrix Nuclear-Norm, we conducted a series of ablation studies. Due to space constraints, detailed results are provided in A.1 (appendix) to maintain brevity in the main text. These experiments included evaluations across different model families, such as Cerebras-GPT and Pythia, as well as comparisons of various data sampling strategies.The results demonstrate that the Matrix Nuclear-Norm consistently performs well across different model scales and sampling variations. This not only confirms its applicability across diverse models but also verifies its stability and reliability in handling large-scale datasets. We also provide an ablation study in the appendix, further proving the method’s efficiency and accuracy in evaluating LLMs.
@@ 8 Limitations
Although Matrix Nuclear-Norm performs well in evaluating the performance of LLMs, it still has some limitations. First, since Matrix Nuclear-Norm’s computation relies on the model’s hidden states, the evaluation results are sensitive to both the model architecture and the training process. As a result, under different model designs or training settings, especially for models like GPT-1.3B and GPT-2.7B, inconsistencies in Matrix Nuclear-Norm’s performance may arise, limiting its applicability across a wider range of models. Additionally, while Matrix Nuclear-Norm offers computational efficiency advantages over traditional methods, it may still face challenges with resource consumption when evaluating extremely large models. As model sizes continue to grow, further optimization of Matrix Nuclear-Norm’s computational efficiency and evaluation stability is required.
@@ A.1.2 Sampling Strategy
Table 6: Ablation study of differnet sampling strategies on the Wikimedia[18] dataset.
@@
A.2 Supplementary Experiment Results
@@
    Tables A.2 and A.2 present the Matrix Nuclear-Norm evaluation results during the inference process for Llama-3 and QWEN-2.
@@
    Table 11 contains the complete results for the comparison of Matrix Nuclear-Norm and other metrics based on Cerebras-GPT family considered in Figure 2(d).
@@
Figure 6: Pythia: Time Comparison of Matrix Entropy and Nuclear-Norm
Model Size Matrix Entropy Time (s) Matrix Nuclear-Norm Time (s) Ratio
14M 52.8669 22.2652 2.3772
31M 114.0820 28.1842 4.0477
70M 320.6641 24.3188 13.1855
160M 631.9762 41.6187 15.1817
410M 1040.9764 80.9814 12.8481
1B 4650.1264 114.0639 40.8387
1.4B 6387.0392 347.8670 18.3858
2.8B 8127.1343 343.3888 23.6778
6.9B 28197.8172 816.6332 34.5350
12B 47273.5235 1276.1128 37.0485
Table 7: Pythia Model: Matrix Entropy vs. Matrix Nuclear-Norm Time Comparison
@@
Table 9: Matrix Nuclear-Norm in Llama3 esponses
@@
Table 10: Language modeling indicators on openbookqa, winogrande and piqa.
@@
Table 11: Language modeling indicators on the Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets.
Table 12: The language modeling indicators (where lower values indicate better performance) for the Pythia models were evaluated on the Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets.
@@
Table 13: The prompts selected from OpenOrca[30] dataset.
@@
A.3 Analysis of Algorithmic Complexity

### v3

@@ 5.2 Matrix Nuclear-Norm Observation
5.2.1 Comparing Computational Time
@@
The significant efficiency gain is due to the lower complexity of Matrix Nuclear-Norm, $O(m\cdot n+n\log n)$, versus Matrix Entropy’s $O(n^{3})$, where $m$ is the embedding dimension (columns). This makes it an efficient metric for LLM evaluation, especially for large-scale models.
@@
Model Size ME Time(s) MNN Time(s) Ratio
 [^1] M 623.5 72.7 8.6
256M 1213.1 110.8 10.9
590M 2959.7 184.8 16.0
1.3B 6760.2 379.0 17.8
2.7B 12083.7 732.6 16.5
6.7B 38791.2 1598.4 24.3
13B 59028.4 2984.2 19.8
Table 1: Cerebras-GPT: Time Comparison between Matrix Entropy (ME) and Matrix Nuclear-Norm (MNN)
@@ 5.2.2 Scaling Law of Matrix Nuclear-Norm
To affirm Matrix Nuclear-Norm’s efficacy as an evaluative metric, we evaluated Cerebras-GPT models on four datasets including dolly-15k, Wikipedia, openwebtext2, and hh-rlhf comparing Matrix Nuclear-Norm, matrix entropy, perplexity, and loss. As shown in Table 10, Matrix Nuclear-Norm decreases consistently with model size, indicating better data compression and processing in larger models. This trend (Figure 2(b)) validates Matrix Nuclear-Norm’s utility across datasets. Notably, anomalies at the 2.7B and 13B highlight areas needing further exploration.
(a) Matrix Entropy
(b) Matrix Nuclear-Norm
Figure 2: Comparison of Matrix Nuclear-Norm, matrix entropy when model scales up.
@@ 5.3.1 Sentence Operation Experiments
Figure 3 shows sentence manipulations impact Matrix Nuclear-Norm values. These values decrease with model size, in line with established scaling laws similar to those governing matrix entropy and perplexity (PPL). As models grow larger, they can capture data patterns more efficiently, reducing redundant information representation, which directly lowers the nuclear norm.
The ranking Reverse $>$ Shuffle & Reverse $>$ Shuffle $>$ Base reflects input disruption. Reverse flips the sentence, introducing maximum disorder and causing a large norm increase. Shuffle only partially rearranges elements, leading to a smaller rise. The unaltered Base condition enables optimal compression.
Notably, the 2.7B model has slightly higher Shuffle and Base values than the 1.3B model, yet this doesn’t challenge the conclusion that larger models compress better. The norm increases with text length because longer texts carry more information, increasing entropy and computational complexity. More data means more potential redundancy for the model to process, driving up the norm value. These results clarify model behavior in relation to size, input structure, and length.
@@ 5.3.2 Analysis of Length Dynamics
Figure 4: The Matrix Nuclear-Norm values increase consistently with longer text input lengths, reflecting the model’s ability to capture more information.
The analysis reveals that Matrix Nuclear-Norm generally increase as input length rises, aligning with our expectations (see Figure 4). Longer inputs necessitate that the model manage and compress more information, which naturally leads to higher Matrix Nuclear-Norm. Most models exhibit this trend, indicating effective handling of the increased information load.
However, the Cerebras-GPT-2.7B and Cerebras-GPT-13B models display anomalies in their Matrix Nuclear-Norm values at 64 and 128 tokens, where the value at 128 tokens is lower than that at 64 tokens. This discrepancy may be attributed to these models employing different information compression mechanisms or optimization strategies tailored to specific input lengths, allowing for more effective compression at those lengths.
Overall, aside from a few outliers, the results largely conform to expectations, demonstrating that Matrix Nuclear-Norm values increase with input length, reflecting the greater volume and complexity of information that models must handle.To address the observed trend of rising Matrix Nuclear-Norm values with longer sentences, we incorporated a normalization step in our methodology via dividing the Matrix Nuclear-Norm values by the sentence length. This adjustment helps mitigate any biases introduced by models that tend to generate longer sentences during inference.
@@ 5.3.3 Analysis of Prompt Learning
The experimental results (shown in Table 2) indicate that we performed inference on different sizes of Cerebras-GPT models using three carefully selected prompts (shown in Table 12) and calculated the Matrix Nuclear-Norm values of their responses. As the model size increased, the Matrix Nuclear-Norm values gradually decreased, demonstrating that larger models possess greater information compression capabilities. The prompts significantly influenced Matrix Nuclear-Norm, with variations reflecting the models’ responses to prompt complexity. Specifically, Cerebras-GPT-1.3B showed a notable decrease in Matrix Nuclear-Norm after the input prompts, indicating its sensitivity to them, while Cerebras-GPT-2.7B exhibited smaller changes. In contrast, Cerebras-GPT-6.7B displayed minimal variation across all prompts, suggesting stable performance regardless of prompt detail. Overall, more detailed prompts resulted in larger information volumes in the model’s responses, leading to corresponding changes in Matrix Nuclear-Norm values.
Table 2: Results of prompt learning without Prompt and with (Prompt 1, 2, 3) the use of prompts. Incorporating prompts as prefixes before the QA pairs enhances the models’ ability to achieve better compression.
@@
6 Evaluating and Ranking LLMs
@@ 6.1 Inference-Based Model Assessment
For instance, the Llama-3 70B model demonstrated superior compression capabilities compared to its 8B counterpart, as reflected by significantly lower Matrix Nuclear-Norm values across both benchmarks (see Table 7). A similar trend was observed in the Vicuna family, where Matrix Nuclear-Norm values consistently decreased from 0.4623 for the 7B model to 0.3643 for the 33B model on the AlpacaEval dataset, indicating progressive improvements in information handling (see Table 3). Additionally, the DeepSeek models exhibited a consistent decrease in Matrix Nuclear-Norm values as model size increased, further demonstrating the metric’s validity.
@@
Model Data 1.3B 6.7B 7B
DeepSeek Alpaca 0.4882 0.3472 0.3352
Arena 0.5754 0.4175 0.4357
@@
6.2 Matrix Nuclear-Norm for Model Ranking
@@
     MODEL Matrix Nuclear-Norm
@@
QWEN 2-7B 0.5989 0.5751 0.5870
Mistral-7B 0.4980 0.5126 0.5053
QWEN 1.5-7B 0.4866 0.5165 0.5016
LLaMA 2-7B 0.4648 0.5038 0.4843
Vicuna-7B 0.4623 0.4824 0.4724
Gemma-7B 0.3759 0.3998 0.3879
DeepSeek-7B 0.3352 0.4357 0.3855
QWEN 1.5-72B 0.5291 0.5065 0.5178
QWEN 2-72B 0.5261 0.4689 0.4975
Llama 3-70B 0.4935 0.4967 0.4951
Llama 2-70B 0.3862 0.4086 0.3974
Table 4: Descending Competence Rankings via Matrix Nuclear Norm: Small and Large LMs
To validate the design rationale and robustness of the Matrix Nuclear-Norm, we conducted a series of ablation studies. Due to space constraints, detailed results are provided in A.1 (appendix) to maintain brevity in the main text. These experiments included evaluations across different model families, such as Cerebras-GPT and Pythia, as well as comparisons of various data sampling strategies.The results demonstrate that the Matrix Nuclear-Norm consistently performs well across different model scales and sampling variations. This not only confirms its applicability across diverse models but also verifies its stability and reliability in handling large-scale datasets. We also provide an ablation Cerebras-GPT: study in the appendix, further proving the method’s efficiency and accuracy in evaluating LLMs.
@@ 8 Limitations
Although Matrix Nuclear-Norm (MNN) performs well in evaluating LLM performance, it has three main limitations. First, as MNN computation relies on hidden states, the results are sensitive to model architecture and training processes. This may cause performance inconsistencies across different model designs or training settings (particularly between Cerebras-GPT-1.3B and Cerebras-GPT-2.7B), potentially limiting broader applicability. Second, while MNN offers computational advantages over traditional methods, it may still face resource challenges when evaluating extremely large models, requiring further optimization for scalability.
Third, our current implementation uses MNN primarily as an evaluation metric rather than a training objective. However, we recognize its potential for analyzing information compression dynamics during training, which could provide valuable insights into model optimization. Future work should explore this direction while addressing the method’s sensitivity to architectural variations.
Notably, despite observed anomalies in specific configurations, MNN demonstrates consistent computational efficiency and accuracy across various model sizes and data sampling strategies. We will enhance our discussion of these performance variations to better clarify the method’s robustness boundaries and operational constraints. These limitations highlight the need for continued research into architecture-agnostic evaluation frameworks and optimized computation strategies as language models scale.
@@ 9 Ethics Statement
10 Reproducibility
We emphasize the importance of reproducibility in the development and evaluation of our newly proposed metric, Matrix Nuclear-Norm. To facilitate reproducibility, we provide detailed information regarding our data processing and parameter settings:
Data Processing and Parameter Settings: We outline the preprocessing steps applied to each dataset, ensuring that other researchers can accurately replicate our methodology. All hyperparameters and configuration settings used during the experiments are specified in the code, offering clarity on the experimental conditions.
Experimental Procedures: We detail the specific steps required to evaluate the Matrix Nuclear-Norm, including its application to each dataset and the metrics used for performance assessment.
Code Availability: Our implementation code, evaluation scripts, and pretrained models will be made publicly available upon acceptance of this paper, enabling others to reproduce our experiments and validate our findings.
By adhering to these guidelines, we aim to ensure that our work is accessible and reproducible for future research endeavors.
@@ A.1.1 Different Model Family
(a) Cross-Entropy Loss
(b) Perplexity
Figure 5: Comparison of loss, and perplexity when model scales up.
@@ A.1.2 Sampling Strategy
Table 5: Ablation study of different sampling strategies on the Wikimedia(Foundation, 2024) dataset.
@@
A.2 Model Selection and Datasets for Analysis
Model Selection. To investigate language model scaling, we employed a diverse set of transformer-based large language models (LLMs) across varying parameter sizes. A key focus of our analysis was the Cerebras-GPT model (Gao et al., 2020), which ranges from 111 million to 13 billion parameters, providing a comprehensive look at scaling effects in pre-trained models. Additionally, we included scaled versions of the Pythia model (Biderman et al., 2023), with parameter counts ranging from 14 million to 12 billion, enabling a broader analysis of model performance across different scales.
To ensure a well-rounded evaluation, we also tested a variety of models, including the DeepSeek series (1.3B, 6.7B, 7B) (Guo et al., 2024), Llama3 series (8B, 70B) (Dubey et al., 2024), QWEN 2 series (0.5B, 1.5B, 7B, 72B) (Yang et al., 2024), and Vicuna models (7B, 13B, 33B) (Chiang et al., 2023). For additional comparative insights, we included models of similar scale, such as Gemma-7B (Team et al., 2024) and Mistral-7B (Jiang et al., 2023).
Datasets for Analysis. Our experiments were conducted using several key benchmark datasets. We selected AlpacaEval(Dubois et al., 2024) and ChatBot Arena (Zheng et al., 2023) as the primary datasets for model evaluation. Additionally, subsets from Wikipedia (Foundation, 2024) and OpenWebText2 (Skylion007, 2019) were utilized to track variations in Matrix Nuclear-Norm values, especially with the Cerebras-GPT models.
To validate the Matrix Nuclear-Norm metric, we employed the dolly-15k dataset (Conover et al., 2023) for instruction tuning and the hh-rlhf dataset (Bai et al., 2022) for reinforcement learning with human feedback (RLHF). Further evaluations were performed on benchmark datasets such as OpenBookQA (Mihaylov et al., 2018), Winogrande (Sakaguchi et al., 2021), and PIQA (Bisk et al., 2020). Lastly, prompt learning experiments with the OpenOrca dataset (Lian et al., 2023b) provided a comprehensive framework for assessing the Matrix Nuclear-Norm’s effectiveness across a variety of inference tasks.
A.3 Supplementary Experiment Results
@@
    Tables 7 and 6 present the Matrix Nuclear-Norm evaluation results during the inference process for Llama-3 and QWEN-2.
@@
    Table 10 contains the complete results for the comparison of Matrix Nuclear-Norm and other metrics based on Cerebras-GPT family considered in Figure 2(b).
@@
Table 7: Matrix Nuclear-Norm in Llama 3 Responses
Figure 7: Pythia: Time Comparison of Matrix Entropy and Nuclear-Norm
Model Size ME (s) MNN (s) Ratio
14M 52.9 22.3 2.4
31M 114.1 28.2 4.0
70M 320.7 24.3 13.2
160M 632.0 41.6 15.2
410M 1040.9 81.0 12.8
1B 4650.1 114.1 40.8
1.4B 6387.0 347.9 18.4
2.8B 8127.1 343.4 23.7
6.9B 28197.8 816.6 34.5
12B 47273.5 1276.1 37.0
Table 8: Pythia Model: Matrix Entropy(ME) vs. Matrix Nuclear-Norm(MNN) Time Comparison
@@
Table 9: Language modeling indicators on openbookqa, winogrande and piqa.Except for the matrix nuclear norm, the data is sourced from Wei et al. (2024)
Table 10: The table illustrates the performance metrics for a range of GPT models on the Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets, encompassing matrix entropy, loss, and perplexity. Except for the matrix nuclear norm, the data is sourced from Wei et al. (2024), underscoring the relationship between model scale and its performance.
@@
Table 11: Language modeling indicators for Pythia models across Dolly-15k, Wikipedia, OpenWebText2, and HH-RLHF datasets (lower values indicate better performance). Except for the matrix nuclear norm, data is derived from Wei et al. (2024), showcasing the correlation between model scale and performance.
@@
Table 12: The prompts selected from OpenOrca(Lian et al., 2023b) dataset.
GPT MODEL SIZE
LENGTH  [^1] M 256M 590M 1.3B 2.7B 6.7B 13B
64 0.4574 0.4125 0.3787 0.3486 0.4053 0.3315 0.4148
128 0.5293 0.4680 0.4270 0.3835 0.4143 0.3477 0.4032
512 0.7883 0.6978 0.6251 0.5554 0.5265 0.4468 0.4422
1024 0.9132 0.8787 0.7802 0.6953 0.6351 0.5383 0.5028
Table 13: Analysis of Length Dynamics
A.4 Analysis of Algorithmic Complexity